# On silting complexes associated to $n$-silting modules

### v1

@@ Introduction
In [25, 4.8], it was asked if, for any $n$-silting module with respect to a projective $n$-presentation $\Sigma$, the complex $\Sigma$ constitutes an $(n+1)$-term silting complex, generalizing the known bijection from the $n=1$ case. We give a partial answer to this in Theorem 2.8, by first showing that any $(n+1)$-term silting complex $\Sigma$ with vanishing intermediate cohomology (i.e., $H^{i}(\Sigma)=0$ for all n<i<0-n<i<0) is a projective $n$-presentation of an $n$-silting module. Furthermore, we show in Theorems 2.8 and 2.9 that the converse is also true if $A$ satisfies certain condition, which is always satisfied in case $A$ is left hereditary or commutative noetherian. We also establish the dual claim for $n$-cosilting modules. We also give a characterization in Lemma 2.5 of the situation in which an $n$-silting module gives rise to a silting complex in terms of certain condition on the kernel of the induced cotorsion pair in a suitable category of representations over the base ring, based on the work of Marks and Vitória [27].
Finally, we specialize to the setting of Commutative Algebra, where the base ring $A$ is a commutative noetherian ring, and provide more substantial motivation for studying nn-(co)silting modules by both showing that the corresponding (co)silting complexes enjoy certain nice properties and by providing explicit constructions of nn-(co)silting modules which are neither nn-(co)tilting modules nor 11-(co)silting modules. More precisely, we show in Theorem 3.2 that the (co)silting complex corresponding to any nn-(co)silting module is, in this setting, always a (co)tilting complex, that is, the associated t-structure is of derived type. This generalizes the previously known case for $n=1$ of Pavon and Vitória [31, Corollary 5.12], see also [23, Theorem 6.18] to arbitrary $n$. Note that, in general, determining whether a silting complex is a tilting complex is a subtle problem, as demonstrated in this setting e.g. in [23, §7]. We then proceed to construct the promised examples over local rings which fail to be Cohen–Macaulay, but enjoy vanishing of intermediate local cohomology, these were classically constructed by Sharp [34]. In particular, the cosilting complex constructed in Example 3.8 is neither a cotilting module nor is the induced t-structure restrictable in the sense of [31], and thus it demonstrates Theorem 3.2 as a new way to establish the derived type.
@@ 1.2. $n$-tilting modules and $n$-cotilting modules
If an $n$-tilting module $T$ satisfies the following stronger version of $(T1)$: $T$ has a projective resolution $0\rightarrow P^{n}\rightarrow\cdots\rightarrow P^{0}\rightarrow T\rightarrow 0$ with each $P^{i}$ finitely generated, then $T$ is called a classical $n$-tilting module. By [7, Theorem 3.11], $T$ is $n$-tilting if and only if $\mathrm{Pres}^{n-1}(T)=T^{\perp_{\infty}}$.
@@
By [7, Theorem 3.11], $C$ is $n$-cotilting if and only if $\mathrm{Copres}^{n-1}(C)={{}^{\perp_{\infty}}C}$.
@@ 1.3. Silting and cosilting complexes
Recall from [33] that any silting complex $X$ comes attached with a t-structure in $\mathbf{D}(A)$ of the form $(X^{\perp_{>0}},X^{\perp_{\leq 0}})$, called the associated silting t-structure. We say that two silting complexes are equivalent if they give rise to the same t-structure. It turns out that two silting complexes $X,X^{\prime}$ are equivalent precisely if $\mathrm{Add}(X)=\mathrm{Add}(X^{\prime})$, see [33, Lemma 4.5]. Similarly, any cosilting complex gives rise to a cosilting t-structure $({}^{\perp_{>0}}X,{}^{\perp_{\leq 0}}X)$, leading to the dual notion of equivalence of cosilting complexes.
@@
[10, Theorem 3.11] A complex $X\in\mathbf{K}^{b}(\mathrm{Proj}(A))$ is silting if and only if the following two conditions hold:
@@ 1.4. Silting complexes as tilting modules
We mark down a crucial property of this identification, see [27, Lemma 2.4]. Let $M\in\mathrm{Mod}(AQ_{n+1}/I)$, $P\in\mathrm{Mod}_{\mathcal{P}}(AQ_{n+1}/I)$, and $I\in\mathrm{Mod}_{\mathcal{I}}(AQ_{n+1}/I)$, then we have for any $i>0$ the isomorphisms:
@@
By [27, Theorem 2.10], we obtain the following bijections.
@@ 1.5. $n$-silting modules and $n$-cosilting modules
Suppose that $\Sigma$ is a projective $n$-presentation of $T$, so that $T=H^{0}(\Sigma)$. Recall from [25, Definition 3.1] that $T$ is called an $n$-silting module (with respect to $\Sigma$) if $\mathrm{Pres}^{n-1}(T)=\mathcal{D}_{\Sigma}$. In this case, the class $\mathrm{Pres}^{n-1}(T)$ is called an $n$-silting class.
@@
Suppose that $\Omega$ is an injective $n$-copresentation of $C$, so that $C=H^{0}(\Omega)$. Recall from [25, Definition 4.1] that $C$ is called an $n$-cosilting module (with respect to $\Omega$) if $\mathrm{Copres}^{n-1}(C)=\mathcal{B}_{\Omega}$. In this case, the class $\mathrm{Copres}^{n-1}(C)$ is called an $n$-cosilting class.
@@
$(1)$ A $1$-silting module is just the silting module in the sense of [4]. A 1-cosilting module is just the cosilting module in the sense of [11]. From [4, Theorem 4.9], we know that for an arbitrary ring there is a bijection between (not necessarily finitely generated) silting modules and (not necessarily compact) 2-silting complexes.
$(2)$ Let $T$ be an $A$-module. By [25, Propositions 3.3 and 4.3], $T$ is $n$-tilting if and only if $T$ is a $n$-silting AA-module with respect to a projective n-presentation $P_{n}\stackrel{{\scriptstyle\sigma}}{{\rightarrow}}P_{n-1}\rightarrow\cdots\rightarrow P_{0}$ with $\sigma$ monic. In addition, $C$ is $n$-cotilting if and only if $C$ is a $n$-cosilting module with respect to an injective $n$-copresentation $E_{0}\rightarrow\cdots\rightarrow E_{n-1}\stackrel{{\scriptstyle\omega}}{{\rightarrow}}E_{n}$ with $\omega$ epic.
@@ 2. $n$-silting modules and silting complexes
The following lemma is inspired by [26, Lemma 5.4].
@@
We know from [27, Lemma 2.5(6)] that the projective dimension of $\Sigma$ is at most $n$ in $\mathrm{Mod}(AQ_{n+1}/I)$. By [15, §6], there exists a cotorsion pair $(^{\perp_{\infty}}(\Sigma^{\perp_{\infty}}),\Sigma^{\perp_{\infty}})$ in $\mathrm{Mod}(AQ_{n+1}/I)$ induced by $\Sigma$. Using condition $({\dagger})$, we obtain that the class
@@ 3. $(iii)$
Let $E=\bigoplus^{n}_{i=1}\overline{A}_{j}$. Since $E$ is projective and injective in $\mathrm{Mod}_{\mathcal{P}}(AQ_{n+1}/I)$ by [27, Lemma 2.5], the object $Z:={\Sigma}\oplus E$ is also partial $n$-tilting in $\mathrm{Mod}(AQ_{n+1}/I)$ and $Z^{\perp_{\infty}}=\Sigma^{\perp_{\infty}}$. Since $E$ is contractible when viewed as a cochain complex, we have $\Phi(Z)\cong\Phi(\Sigma)$ in $\mathbf{D}(A)$, and thus it is sufficient to establish the condition $(iii)$ for $Z$ instead of for $\Sigma$.
@@
Finally, let $P$ be the $\mathrm{Mod}(AQ_{n+1}/I)$-module corresponding to the stalk cochain complex $(\cdots\to 0\to 0\to A\to 0\to 0\to\cdots)$ concentrated in degree 0. Then $P$ is projective as an $\mathrm{Mod}(AQ_{n+1}/I)$-module by [27, Lemma 2.4], and thus the preceding paragraph applies to it, showing that $A$ is in the smallest localizing subcategory of $\mathbf{D}(A)$ generated by $Z$. As $A^{\perp_{\mathbb{Z}}}=0$, the condition $(iii)$ follows.
@@
(1) If $A$ is left hereditary then any object of $\mathbf{D}(A)$ is isomorphic to a direct sum of shifts of its cohomology modules, see e.g. [24, §1.6], from which the first claim immediatelly follows.
@@
By Theorem 2.8, we know that $\Omega$ is a cosilting complex and so is pure-injective as an object of $\mathbf{D}(A)$ by [27, Proposition 3.10]. Then $C=H^{0}(\Omega)$ is pure-injective by [32, Theorem 17.3.19]. ∎
@@ 3.1. Commutative Algebra setup
The main bijection follows from the classification result [3, Theorem 3.8], see also the exposition in [38, §3] and [23, §3]. In particular, the statement about cofinite type was proved in [22]. The first bullet claim follows from the proof of [3, Theorem 3.8], while the second is discussed in [23, Remark 5.11]. ∎
The following provides a natural generalization of [31, Corollary 5.12], which constitutes the $n=1$ case.
@@
Clearly, the indermediate cohomology of the cosilting complex $\Sigma^{+}$ vanishes and thus $\Sigma^{+}$ represents an $n$-cosilting AA-module. It follows that $\Sigma^{+}$ is a cotilting complex by Theorem 3.2. Then $\Sigma$ is tilting by [20, Corollary 3.5]. ∎
@@
In fact, [20, Theorem 3.4] shows more: If $\Sigma$ is a silting complex over $A$ which represents an $n$-silting module then $\Sigma$ is a decent tilting complex. This means that the heart of the silting t-structure is equivalent to the category of contramodules over the endomorphism ring of $\Sigma$ endowed with certain natural linear topology. Note that this category of contramodules, as well as the dual-analogous category of discrete modules, is then derived equivalent to $\mathrm{Mod}(A)$. See [20] for details.
@@ 3.2. Examples over local rings with vanishing intermediate cohomology
By a construction due to Sharp [34, Proposition 1.4], such rings exist for any choice of non-negative integers $s\leq d$. An explicit example for $s=0$ and $d=2$ is $A=k[[x,y,z]]/(x^{2},xy,xz)$. We will be particularly interested in the case $s=0$, that is, the case of $A$ being of zero depth. Note that in this case, in view of Theorem 3.1, the only $n$-tilting module (resp. $n$-cotilting module) is, up to equivalence, the projective generator (resp. the injective cogenerator). However, we shall show that $n$-silting modules are much more abundant over such $A$. Furthermore, such $A$ is what is called a generalized Cohen–Macaulay local ring, which means by definition that $H^{i}_{\mathfrak{m}}(A)$ is of finite length for all $i<d$. This is because $H^{0}_{\mathfrak{m}}(A)$ is always finite length, being a submodule of $A$, and the intermediate local cohomology modules $H^{i}_{\mathfrak{m}}(A)$ for $0<i<d$ vanish by the assumption. It follows by [19, Theorem 37.4] that $A$ is Cohen–Macaulay on the punctured spectrum, that is, $A_{\mathfrak{p}}$ is Cohen–Macaulay for any non-maximal prime ideal $\mathfrak{p}$.
@@
be the corresponding silting complex as constructed in [23, §4]. We claim that $\Sigma$ represents a $d$-silting module. For any non-maximal prime ideal $\mathfrak{p}$, the local cohomology $\mathbf{R}\Gamma_{V(\mathfrak{p})}(A_{\mathfrak{p}})[\operatorname{ht}(\mathfrak{p})]$ is quasi-isomorphic to a module in degree zero, this reflects $A_{\mathfrak{p}}$ being Cohen–Macaulay, see above. Then $H^{i}(\Sigma)=0$ for all d<i<0-d<i<0. Since $\Sigma$ is a $(d+1)$-term silting complex by [23, §4], we see that $\Sigma$ represents a $d$-silting module. Since the monotone perversity, the function $\operatorname{ht}$ is not constant zero and thus is not bounded by depth, so that $\Sigma$ does not represent an $n$-tilting module for any $n\geq 0$. For any choice $d>1$, the monotone perversity $\operatorname{ht}$ exceeds $1$ in value, and thus $\Sigma$ is not a $2$-term silting complex, and thus does not represent a $1$-silting module.
As per [23, §5], note that the dual cosilting complex $\Sigma^{+}$ is equivalent to $\prod_{\mathfrak{p}\in\operatorname{Spec}(A)}D_{\widehat{R_{\mathfrak{p}}}}$, where $D_{\widehat{R_{\mathfrak{p}}}}$ is the dualizing complex over the $\mathfrak{p}R_{\mathfrak{p}}$-adic completion of the local ring $R_{\mathfrak{p}}$. Since for any non-maximal prime $\mathfrak{p}$ the local ring $R_{\mathfrak{p}}$ is Cohen–Macaulay, $D_{\widehat{R_{\mathfrak{p}}}}$ is isomorphic to the canonical module over the local ring $\widehat{R_{\mathfrak{p}}}$.

### v2

@@ Introduction
In [26, 4.8], it was asked if, for any $n$-silting module with respect to a projective $n$-presentation $\Sigma$, the complex $\Sigma$ constitutes an $(n+1)$-term silting complex, generalizing the known bijection from the $n=1$ case. We give a partial answer to this in Theorem 2.8, by first showing that any $(n+1)$-term silting complex $\Sigma$ with vanishing intermediate cohomology (i.e., $H^{i}(\Sigma)=0$ for all n<i<0-n<i<0) is a projective $n$-presentation of an $n$-silting module. Furthermore, we show in Theorems 2.8 and 2.9 that the converse is also true if $A$ satisfies certain condition, which is always satisfied in case $A$ is left hereditary or commutative noetherian. We also establish the dual claim for $n$-cosilting modules. We also give a characterization in Lemma 2.5 of the situation in which an $n$-silting module gives rise to a silting complex in terms of certain condition on the kernel of the induced cotorsion pair in a suitable category of representations over the base ring, based on the work of Marks and Vitória [28].
Finally, we specialize to the setting of Commutative Algebra, where the base ring $A$ is a commutative noetherian ring, and provide more substantial motivation for studying nn-(co)silting modules by both showing that the corresponding (co)silting complexes enjoy certain nice properties and by providing explicit constructions of nn-(co)silting modules which are neither nn-(co)tilting modules nor 11-(co)silting modules. More precisely, we show in Theorem 3.2 that the (co)silting complex corresponding to any nn-(co)silting module is, in this setting, always a (co)tilting complex, that is, the associated t-structure is of derived type. This generalizes the previously known case for $n=1$ of Pavon and Vitória [32, Corollary 5.12], see also [24, Theorem 6.18] to arbitrary $n$. Note that, in general, determining whether a silting complex is a tilting complex is a subtle problem, as demonstrated in this setting e.g. in [24, §7]. We then proceed to construct the promised examples over local rings which fail to be Cohen–Macaulay, but enjoy vanishing of intermediate local cohomology, these were classically constructed by Sharp [35]. In particular, the cosilting complex constructed in Example 3.8 is neither a cotilting module nor is the induced t-structure restrictable in the sense of [32], and thus it demonstrates Theorem 3.2 as a new way to establish the derived type.
@@ 1.2. $n$-tilting modules and $n$-cotilting modules
If an $n$-tilting module $T$ satisfies the following stronger version of $(T1)$: $T$ has a projective resolution $0\rightarrow P^{n}\rightarrow\cdots\rightarrow P^{0}\rightarrow T\rightarrow 0$ with each $P^{i}$ finitely generated, then $T$ is called a classical $n$-tilting module. By [8, Theorem 3.11], $T$ is $n$-tilting if and only if $\mathrm{Pres}^{n-1}(T)=T^{\perp_{\infty}}$.
@@
By [8, Theorem 3.11], $C$ is $n$-cotilting if and only if $\mathrm{Copres}^{n-1}(C)={{}^{\perp_{\infty}}C}$.
@@ 1.3. Silting and cosilting complexes
Recall from [34] that any silting complex $X$ comes attached with a t-structure in $\mathbf{D}(A)$ of the form $(X^{\perp_{>0}},X^{\perp_{\leq 0}})$, called the associated silting t-structure. We say that two silting complexes are equivalent if they give rise to the same t-structure. It turns out that two silting complexes $X,X^{\prime}$ are equivalent precisely if $\mathrm{Add}(X)=\mathrm{Add}(X^{\prime})$, see [34, Lemma 4.5]. Similarly, any cosilting complex gives rise to a cosilting t-structure $({}^{\perp_{>0}}X,{}^{\perp_{\leq 0}}X)$, leading to the dual notion of equivalence of cosilting complexes.
@@
[11, Theorem 3.11] A complex $X\in\mathbf{K}^{b}(\mathrm{Proj}(A))$ is silting if and only if the following two conditions hold:
@@ 1.4. Silting complexes as tilting modules
We mark down a crucial property of this identification, see [28, Lemma 2.4]. Let $M\in\mathrm{Mod}(AQ_{n+1}/I)$, $P\in\mathrm{Mod}_{\mathcal{P}}(AQ_{n+1}/I)$, and $I\in\mathrm{Mod}_{\mathcal{I}}(AQ_{n+1}/I)$, then we have for any $i>0$ the isomorphisms:
@@
By [28, Theorem 2.10], we obtain the following bijections.
@@ 1.5. $n$-silting modules and $n$-cosilting modules
Suppose that $\Sigma$ is a projective $n$-presentation of $T$, so that $T=H^{0}(\Sigma)$. Recall from [26, Definition 3.1] that $T$ is called an $n$-silting module (with respect to $\Sigma$) if $\mathrm{Pres}^{n-1}(T)=\mathcal{D}_{\Sigma}$. In this case, the class $\mathrm{Pres}^{n-1}(T)$ is called an $n$-silting class.
@@
Suppose that $\Omega$ is an injective $n$-copresentation of $C$, so that $C=H^{0}(\Omega)$. Recall from [26, Definition 4.1] that $C$ is called an $n$-cosilting module (with respect to $\Omega$) if $\mathrm{Copres}^{n-1}(C)=\mathcal{B}_{\Omega}$. In this case, the class $\mathrm{Copres}^{n-1}(C)$ is called an $n$-cosilting class.
@@
$(1)$ A $1$-silting module is just the silting module in the sense of [5]. A 1-cosilting module is just the cosilting module in the sense of [12]. From [5, Theorem 4.9], we know that for an arbitrary ring there is a bijection between (not necessarily finitely generated) silting modules and (not necessarily compact) 2-silting complexes.
$(2)$ Let $T$ be an $A$-module. By [26, Propositions 3.3 and 4.3], $T$ is $n$-tilting if and only if $T$ is a $n$-silting AA-module with respect to a projective n-presentation $P_{n}\stackrel{{\scriptstyle\sigma}}{{\rightarrow}}P_{n-1}\rightarrow\cdots\rightarrow P_{0}$ with $\sigma$ monic. In addition, $C$ is $n$-cotilting if and only if $C$ is a $n$-cosilting module with respect to an injective $n$-copresentation $E_{0}\rightarrow\cdots\rightarrow E_{n-1}\stackrel{{\scriptstyle\omega}}{{\rightarrow}}E_{n}$ with $\omega$ epic.
@@ 2. $n$-silting modules and silting complexes
The following lemma is inspired by [27, Lemma 5.4].
@@
We know from [28, Lemma 2.5(6)] that the projective dimension of $\Sigma$ is at most $n$ in $\mathrm{Mod}(AQ_{n+1}/I)$. By [16, §6], there exists a cotorsion pair $(^{\perp_{\infty}}(\Sigma^{\perp_{\infty}}),\Sigma^{\perp_{\infty}})$ in $\mathrm{Mod}(AQ_{n+1}/I)$ induced by $\Sigma$. Using condition $({\dagger})$, we obtain that the class
@@ 3. $(iii)$
Let $E=\bigoplus^{n}_{i=1}\overline{A}_{j}$. Since $E$ is projective and injective in $\mathrm{Mod}_{\mathcal{P}}(AQ_{n+1}/I)$ by [28, Lemma 2.5], the object $Z:={\Sigma}\oplus E$ is also partial $n$-tilting in $\mathrm{Mod}(AQ_{n+1}/I)$ and $Z^{\perp_{\infty}}=\Sigma^{\perp_{\infty}}$. Since $E$ is contractible when viewed as a cochain complex, we have $\Phi(Z)\cong\Phi(\Sigma)$ in $\mathbf{D}(A)$, and thus it is sufficient to establish the condition $(iii)$ for $Z$ instead of for $\Sigma$.
@@
Finally, let $P$ be the $\mathrm{Mod}(AQ_{n+1}/I)$-module corresponding to the stalk cochain complex $(\cdots\to 0\to 0\to A\to 0\to 0\to\cdots)$ concentrated in degree 0. Then $P$ is projective as an $\mathrm{Mod}(AQ_{n+1}/I)$-module by [28, Lemma 2.4], and thus the preceding paragraph applies to it, showing that $A$ is in the smallest localizing subcategory of $\mathbf{D}(A)$ generated by $Z$. As $A^{\perp_{\mathbb{Z}}}=0$, the condition $(iii)$ follows.
@@
(1) If $A$ is left hereditary then any object of $\mathbf{D}(A)$ is isomorphic to a direct sum of shifts of its cohomology modules, see e.g. [25, §1.6], from which the first claim immediatelly follows.
@@
By Theorem 2.8, we know that $\Omega$ is a cosilting complex and so is pure-injective as an object of $\mathbf{D}(A)$ by [28, Proposition 3.10]. Then $C=H^{0}(\Omega)$ is pure-injective by [33, Theorem 17.3.19]. ∎
@@ 3.1. Commutative Algebra setup
The main bijection follows from the classification result [4, Theorem 3.8], see also the exposition in [38, §3] and [24, §3]. In particular, the statement about cofinite type was proved in [23]. The first bullet claim follows from the proof of [4, Theorem 3.8], while the second is discussed in [24, Remark 5.11]. ∎
The following provides a natural generalization of [32, Corollary 5.12], which constitutes the $n=1$ case.
@@
Clearly, the indermediate cohomology of the cosilting complex $\Sigma^{+}$ vanishes and thus $\Sigma^{+}$ represents an $n$-cosilting AA-module. It follows that $\Sigma^{+}$ is a cotilting complex by Theorem 3.2. Then $\Sigma$ is tilting by [21, Corollary 3.5]. ∎
@@
In fact, [21, Theorem 3.4] shows more: If $\Sigma$ is a silting complex over $A$ which represents an $n$-silting module then $\Sigma$ is a decent tilting complex. This means that the heart of the silting t-structure is equivalent to the category of contramodules over the endomorphism ring of $\Sigma$ endowed with certain natural linear topology. Note that this category of contramodules, as well as the dual-analogous category of discrete modules, is then derived equivalent to $\mathrm{Mod}(A)$. See [21] for details.
@@ 3.2. Examples over local rings with vanishing intermediate cohomology
By a construction due to Sharp [35, Proposition 1.4], such rings exist for any choice of non-negative integers $s\leq d$. An explicit example for $s=0$ and $d=2$ is $A=k[[x,y,z]]/(x^{2},xy,xz)$. We will be particularly interested in the case $s=0$, that is, the case of $A$ being of zero depth. Note that in this case, in view of Theorem 3.1, the only $n$-tilting module (resp. $n$-cotilting module) is, up to equivalence, the projective generator (resp. the injective cogenerator). However, we shall show that $n$-silting modules are much more abundant over such $A$. Furthermore, such $A$ is what is called a generalized Cohen–Macaulay local ring, which means by definition that $H^{i}_{\mathfrak{m}}(A)$ is of finite length for all $i<d$. This is because $H^{0}_{\mathfrak{m}}(A)$ is always finite length, being a submodule of $A$, and the intermediate local cohomology modules $H^{i}_{\mathfrak{m}}(A)$ for $0<i<d$ vanish by the assumption. It follows by [20, Theorem 37.4] that $A$ is Cohen–Macaulay on the punctured spectrum, that is, $A_{\mathfrak{p}}$ is Cohen–Macaulay for any non-maximal prime ideal $\mathfrak{p}$.
@@
be the corresponding silting complex as constructed in [24, §4]. We claim that $\Sigma$ represents a $d$-silting module. For any non-maximal prime ideal $\mathfrak{p}$, the local cohomology $\mathbf{R}\Gamma_{V(\mathfrak{p})}(A_{\mathfrak{p}})[\operatorname{ht}(\mathfrak{p})]$ is quasi-isomorphic to a module in degree zero, this reflects $A_{\mathfrak{p}}$ being Cohen–Macaulay, see above. Then $H^{i}(\Sigma)=0$ for all d<i<0-d<i<0. Since $\Sigma$ is a $(d+1)$-term silting complex by [24, §4], we see that $\Sigma$ represents a $d$-silting module. Since the monotone perversity, the function $\operatorname{ht}$ is not constant zero and thus is not bounded by depth, so that $\Sigma$ does not represent an $n$-tilting module for any $n\geq 0$. For any choice $d>1$, the monotone perversity $\operatorname{ht}$ exceeds $1$ in value, and thus $\Sigma$ is not a $2$-term silting complex, and thus does not represent a $1$-silting module.
As per [24, §5], note that the dual cosilting complex $\Sigma^{+}$ is equivalent to $\prod_{\mathfrak{p}\in\operatorname{Spec}(A)}D_{\widehat{R_{\mathfrak{p}}}}$, where $D_{\widehat{R_{\mathfrak{p}}}}$ is the dualizing complex over the $\mathfrak{p}R_{\mathfrak{p}}$-adic completion of the local ring $R_{\mathfrak{p}}$. Since for any non-maximal prime $\mathfrak{p}$ the local ring $R_{\mathfrak{p}}$ is Cohen–Macaulay, $D_{\widehat{R_{\mathfrak{p}}}}$ is isomorphic to the canonical module over the local ring $\widehat{R_{\mathfrak{p}}}$.